OpenStudy (anonymous):

Addition method 4y+5z=-2 5y-3z=16 Multiplied 1st with 3 and 2nd with 5 12y+15z=-6 and 25y-15z=16 37y= 10 divided by 37 ended up with y=10/37 Am I heading down the right path or?

hero (hero):

Well let me use a different method and see if I come up with the same thing as you.

hero (hero):

5y-3z=16 4y+5z=-2 Subtract the second equation from the first yields: y - 8z = 18 y = 8z + 18 Substituting y in the original system: 5(8z + 18) - 3z = 16 40z + 90 - 3z = 16 37z = -74 z = -74/37 `z = -2` y = 8(-2) + 18 y = -16 + 18 `y = 2`

OpenStudy (anonymous):

Thanks

hero (hero):

yw